# DietGlance: Dietary Monitoring and Personalized Analysis at a Glance with Knowledge-Empowered AI Assistant

### v1

@@ 5.3.1 Evaluation Settings
Recognizing the variability in nutrient estimation, we recruited ten food and nutrition experts, including six nutrition interns, three researchers, and one Accredited Practicing Dietitian. They independently analyzed meal data (the ground truth meal images across the eating process), referencing the database provided by the Centre of Food Safety of [masked for anonymous compliance] database, with each meal randomly assigned to five experts for estimation. Each expert received about \$25 for compensation.
@@ 5.3.2 Evaluation Metrics
We measured expert agreement using the Intraclass Correlation Coefficient (ICC), showing high reliability of the average estimation from experts with ICC(2,k) > 0.9 (more details in the supplementary material 4). Therefore, to evaluate DietGlance’s performance, we compared its estimation to the expert averages using Mean Absolute Percentage Error (MAPE) due to varying nutrient units and scales, calculated as: ${MAPE=\frac{1}{N}\sum_{n=1}^{N}|\frac{\overline{E}_{n}-S_{n}}{\overline{E}_{n}}|\times 100\%}$, where $N$ is the number of meals. $\overline{E}_{n}$ and $S_{n}$ are the average estimations of experts and the system, respectively.
@@ 6 Study II: Four-Week Longitudinal Study
To evaluate the long-term effect of DietGlance on users’ dietary patterns, we recruited 16 participants from Study I for a 4-week longitudinal study with IRB approval, including 4 females and 12 males (P1, P6, P7, P10, P13, P15, P18, P20, P21, exp., P3, P8, P9, P14, P16, P17, P22, inexp.), aged 24 to 33 ($M=28.3$, $SD=2.4$).
Building on the design requirements and findings from Study I, this study aims to explore how exp. and inexp. participants perceive and use DietGlance over time, and how it affects their dietary patterns.
@@ 6.1 Study Settings
  • Incorporated users’ habits (e.g., drinking Coke Zero instead of regular Coke) into prompts to refine diet identification (details in Table A7)
@@
During the study, participants were encouraged to use Aria Glasses covering as many ingestive behaviors (meals and snacks) as possible to gain comprehensive insights into their habits and use the interface to review their dietary monitoring results. The interface was available 24/7, except from 12:00 am to 5:59 am for potential maintenance and backup, and was tracked via Google Analytics. Meal data older than seven days was archived, with analysis based on the most recent seven days to reflect current dietary patterns. Participants provided weekly feedback through questionnaires and interviews (details in Tables A9 and A8). Participants received about \$17 every week and could review their monitoring and analysis results whenever the interface was available.
@@ 6.2.3 Long-Term General Usability, User Satisfaction, and User Experience
Usability improved throughout the study, with SUS scores rising from 76.09 in Study I to 85.00 by the end of Study II, reflecting increased perceived usability. Exp. participants consistently rated the system higher, while inexp. ones showed steady improvement, highlighting adaptability. NPS scores also rose, indicating sustained satisfaction across both groups. User experience, measured through UEQ dimensions, showed positive trends in attractiveness, efficiency, and dependability, while stimulation and novelty declined slightly, reflecting reduced novelty as users became familiar with the system. These findings emphasize the DietGlance’s high usability, satisfaction, and user experience, highlighting the need for tailoring onboarding processes and maintaining engagement, particularly for inexp. users. More details are in the supplementary material 5.
@@ 6.2.4 Long-Term Interaction Behaviors with DietGlance
Participants demonstrated consistent engagement with the system over four weeks, averaging 38.13 usage sessions. Exp. showed slightly higher engagement than inexp. ones. Initial novelty drove higher interactions in Week 1, which gradually stabilized into routine usage by Week 3. The average engagement time per session was 5.95 minutes, with exp. participants spending longer on average, reflecting deeper interaction with the system. The relatively low number of logged updates to diet summaries highlighted the system’s accuracy, while sustained chatbot interactions demonstrated its role in enhancing personalization and feedback. More details are available in the supplementary material 6.
@@ 7 Discussion
In this section, we further discuss the design considerations for future dietary monitoring and analysis systems, followed by a discussion on the generalizability and limitations of this work.
7.1 Design Considerations for Eyewear-based Dietary Monitoring and Analysis Systems
@@ 7.3 Limitations and Future Work
The Aria Glasses, though powerful, are heavy due to unnecessary sensors for our study. Participants reported discomfort from slippage and sweat, especially in summer. Many also required vision correction, but lens replacement was not practical. This discomfort may hinder long-term retention. Future research could explore lighter glasses, attachable sensors for personal eyewear, or alternatives like smart earbuds or watches.
@@
DietGlance accurately identified most foods but struggled with visually similar items. While DietGlance accurately identified most foods, it struggled with visually similar items. Users may need to manually update diet summaries. The system also assumed equal portions in shared meals, potentially leading to discrepancies. Additionally, estimating micronutrients like phosphorus and potassium was challenging due to variability in biological, environmental, and processing factors. Future work could focus on improving micronutrient datasets for higher accuracy.
@@ 8 Conclusion
Appendix A Methodological Transparency & Reproducibility Appendix
Table A1: Class Prompts for Diet Segmentation.
[‘food’, ‘drink’, ‘soup’, ‘rice’, ‘noodle’, ‘meat’, ‘vegetable’, ‘seafood’, ‘fried egg’, ‘cutlet’, ‘curry’, ‘broccoli’, ‘fruit’, ‘fish’, ‘chicken’, ‘lemon’, ‘beef’, ‘pork’, ‘burger’, ‘salad’, ‘cola’, ‘ham’, ‘bread’, ‘cake’, ‘dessert’, ‘sandwich’, ‘bread’, ‘snack’, ‘juice’, ‘sauce’, ‘coffee’, ‘tea’, ‘water’, ‘alcohol’, ‘milk’, ‘beer’, ‘oil’, ‘cup’, ’bottle’, ‘plate’, ‘bowl’, ‘mug’, ‘can’]
Table A2: Diet Identification Prompt.
System: Based on a series of egocentric images taken from the camera on the user’s glasses during a meal session, considering the user information, describe this meal containing the following information: - Names of the food and drink taken by users. Separate the ingredients as detailed as possible. - Based on the images provided, which are taken during the process of eating and drinking, estimate the amount of food (in g) and drink (in ml) at the beginning, and how much percentage had been consumed by the user at the end. Constraints: - If you don’t know the answer, say "I don’t know", don’t try to make up an answer. - Output the answer following the format of the Structured Output Example below. Structured Output Example: - A 10-minute dinner session. The meal includes a bowl of beef noodle soup with various vegetables. - Fried Beef, amount: 80g, consumed percentage: 90% - Lemon Water, amount: 250ml, consumed percentage: 50% User: [Meal Image Log] The provided images in this message were captured on [date] from [time] to [time]. The user is a [age]-year-old [gender], with a weight of [weight] kg and a height of [height] cm. The user has a dietary habit of [habit].
This sentenced was added in the longitudinal study.
Table A3: Nutritional Analysis Prompt.
System: You are an excellent nutritionist, capable of identifying the nutrients contained in food and drinks. Provide a detailed breakdown of nutrients for the given name and amount of food or drink based only on the provided context below. Context: [context] Constraints: - If you don’t know the answer, say "I don’t know", don’t try to make up an answer. - Output the answer following the structured output below. Don’t explain or provide additional information. Don’t add descriptions such as ‘Here are the results’ and ‘Structured Output’. The output should be a plain text list where each line starts with the name of the nutrient, followed by a colon and then the amount and unit of the nutrient. The amount and the unit should be separated by a comma. Structured Output: - energy: amount without unit, kcal - protein: amount without unit, g - carbohydrate: amount without unit, g - total_fat: amount without unit, g - dietary_fibre: amount without unit, g - sugars: amount without unit, g - saturated_fat: amount without unit, g - trans_fat: amount without unit, g - cholesterol: amount without unit, mg - calcium: amount without unit, mg - copper: amount without unit, mg - iron: amount without unit, mg - magnesium: amount without unit, mg - manganese: amount without unit, mg - phosphorus: amount without unit, mg - potassium: amount without unit, mg - sodium: amount without unit, mg - zinc: amount without unit, mg - vitamin_c: amount without unit, mg User: Provide a detailed breakdown of nutrients in [name] with the amount of [amount] [unit]. Present the results following the format of the Structured Output.
Table A4: General and Personalized Dietary Suggestion Generation Prompt.
[…]
Table A5: Nutrition Consumed Evaluation Prompt.
System: You are an excellent nutritionist. You excel in analyzing the nutrient consumed by the user and judge if each nutrient taken is too high, too low, or okay, and provide suggested values for the nutrient based on the user’s information and need. Description of the session: [description of the meal] Nutrition of a meal taken by the user: [nutrient 1: amount, nutrient 2: amount, ...] User information: age [age], gender [gender], weight [weight] kg, height [height] cm, need [need]. Context: [context] Constraints: - If you don’t know the answer, say "not sure", don’t try to make up an answer. - Output one of the three: too high, too low, or okay for each nutrient taken in this eating session, and provide suggested values. The suggested values should follow too high/ too low/ okay, enclosed in parentheses. - Only output the results following the structured output. Don’t explain or provide additional information. - Don’t add descriptions such as ‘here are the results’ and ‘Structured Output.’ Structured Output:
Nutrient: too high/ too low/ okay (suggested value for this meal or snack session, only output the values) User: Analyze each nutrition taken by the user as higher, lower, or okay, and provide suggested values for each nutrient.
Table A6: AI Nutrition Assistant Prompt
System: [chat history]
User: [question]
User: Given the above conversation, generate a search query to look up to get information relevant to the conversation
System: As an excellent nutritionist, you excel in answering users’ questions based on the context provided below and the chat history. Context: [context] Constraints: - If you don’t know the answer, say "I don’t know", don’t try to make up an answer. - Limit the answer to 150 words. MessagesPlaceholder([chat history])
User: [question]. No more than 150 words.
Table A7: Description of participants in Study I, including their unique ID, age, self-declared gender, occupation background, dietary monitoring experience, dietary goal, daily glasses-wearing behaviors, and meal sessions collected. Sixteen participants among them (P1, P3, P6-P10, P13-P18, P20-P22) further attended the Study II.
ID Age Gender Occupation Experience Dietary Goal Glasses Wearing Meals Dietary Habit
P1 28 F Student exp. Lose Weight Never 9 None
P2 28 F Student exp. Lose Weight Never 6 Not Applicable
P3 27 M Student inexp. Gain Muscle All the time 5 None
P4 24 M Student exp. Gain Muscle All the time 5 Not Applicable
P5 30 F Student inexp. keep Fit All the time 4 Not Applicable
P6 28 F Postdoc exp. Mitigate Pimple All the time 7 Never drink coffee or milk tea
P7 26 M Student exp. Lose Weight Never 8 Never eat the bottom bread slice of
hamburgers, only drink sugar-free Coke
P8 29 M Student inexp. Lose Weight Most of the time 6 None
P9 24 F Student inexp. Lose Weight All the time 5 Drink sugar-free coffee at breakfast,
sugar-free milk tea at lunch and dinner
P10 27 M Postdoc exp. None Rarely 5 Drink sugar-free soda
P11 35 M Postdoc exp. Lose Weight Never 3 Not Applicable
P12 60 F Accountant inexp. Lose Weight Never 3 Not Applicable
P13 27 M Student exp. Lose Weight All the time 3 All drinks are sugar free.
P14 31 M Professor inexp. None All the time 2 Drink coffee milk tea at lunch
P15 28 M Student exp. Gain Muscle All the time 3 None
P16 30 F Postdoc inexp. None Never 3 Add sugar to lemon water and lemon tea
P17 32 M Postdoc inexp. Gain Weight All the time 3 Drink hot lemon water
P18 25 M Student exp. Gain Muscle All the time 5 None
P19 29 F Professor inexp. Lose Weight Never 3 Not Applicable
P20 28 M Student exp. Keep Fit All the time 2 Only drink soup or iced lemon tea
P21 33 M Postdoc exp. Lose Fat, All the time 3 Drink sugar-free coke, milk tea,
Gain Muscle and black coffee
P22 30 M Student inexp. None Never 7 None
P23 43 F Professor exp. None Never 7 Not Applicable
P24 46 M Engineer exp. None Never 5 Not Applicable
P25 73 M Retired exp. None Never 4 Not Applicable
P26 26 M Student exp. Lose Weight All the time 3 Not Applicable
P27 25 M Student inexp. Gain Weight All the time 4 Not Applicable
P28 23 M Student exp. None All the time 4 Not Applicable
P29 23 M Student exp. Lose Weight All the time 4 Not Applicable
P30 24 M Student exp. Lose Weight All the time 4 Not Applicable
P31 23 F Student exp. Lose Weight All the time 4 Not Applicable
P32 24 F Student exp. Lose Weight Sometimes 3 Not Applicable
P33 25 F Student inexp. None All the time 2 Not Applicable
‘Not Applicable’ means we did not collect this information for the participant since they did not attend the longitudinal study.
Table A8: The weekly interview script in study II.
Q1. What features did you find most useful or challenging this week? Have you noticed new features you didn’t use in the previous weeks?
Q2. Have you noticed any changes in your dietary habits since using the system this week? If yes, how do you find it? By the information provided by the product? Or your personal subjective sense? Or others?
Q3. Do you have any suggestions for improvements or new features?
Q4. Is there anything else you would like to share about your experience this week?
(Last week) Q5. Is there anything else you would like to share about your experience with the system across the four weeks?
Table A9: The survey scale (Study I and II) and interview questions (Study I). The scale includes System Usability Scale (SUS), Net Promoter Score (NPS), User Experience Questionnaire (UEQ), and questions about users’ perception on the diet identification and logging (R1), nutritional analysis (R2), personalized dietary suggestions (R3), and privacy protection (R4) toward DietGlance.
Scale Item
SUS Q1. I think that I would like to use this system frequently.
Q2. I found the system unnecessarily complex.
Q3. I thought the system was easy to use.
Q4. I think that I would need the support of a technical person to be able to use this system.
Q5. I found the various functions in this system were well integrated.
Q6. I thought there was too much inconsistency in this system.
Q7. I would imagine that most people would learn to use this system very quickly.
Q8. I found the system very cumbersome to use.
Q9. I felt very confident using the system.
Q10. I needed to learn a lot of things before I could get going with this system.
NPS Q11. On a scale from 0 to 10, how likely are you to recommend this system to a friend or colleague?
Q12. What is the primary reason for your score?
UEQ Q13. (attractiveness) The system is attractive.
Q14. (Perspicuity) The system is easy to understand.
Q15. (Efficiency) The system is fast and efficient to use.
Q16. (Dependability) The system is predictable and reliable. (Predicable: users can anticipate how the product will respond to their actions. Reliable: the product consistently performs its functions without errors.)
Q17. (Stimulation) The system is exciting and motivating to use.
Q18. (Novelty) The system is innovative and creative.
Diet Q19. The system is accurate in identifying and logging your meals.
Identification Q20. It is easy to use the system for capturing meals.
and Logging (R1) Q21. Please describe any issues you encountered with the automated logging process
Nutritional Q22. The nutritional analysis provided by the system is accurate
Analysis (R2) Q23. The nutritional analysis is useful for your dietary planning
Q24. Please describe any discrepancies between the system’s nutritional values and your expectations.
Personalized Q25. The dietary suggestions were relevant to your dietary goals.
Dietary Q26. The dietary suggestions were easy to follow.
Suggestions (R3) Q27. I trust the dietary suggestions.
Privacy Q28. I’m confident my personal data is secure with this system.
Protection (R4) Q29. How my data used in this system is transparent.
Q30. Please describe have any concerns about privacy when using the system.
Interview Question
R1 Q1. Describe your overall experience using this system for meal logging.
Q2. Do you think the system is accurate in meal logging?
Q3. Were there any meals or food items that the system struggled to identify?
Q4. What improvements would you suggest for the food identification and logging feature?
R2 Q1. How was your experience with the nutritional analysis feature?
Q2. Were there nutrients or food items where the analysis seemed accurate or inaccurate?
Q3. Do you trust the nutritional analysis results? Why? What elements in the system increase your trust.
Q4. How has the nutritional analysis influenced your dietary decisions?
R3 Q1. How did you use the personalized dietary suggestions in your daily routine?
Q2. Do you think the dietary suggestions are reliable? Why? What elements in the system make them reliable or unreliable?
Q3. Were there any suggestions that were particularly helpful or unhelpful?
Q4. What changes or additional features would you recommend?
R4 Q1. How do you feel about the system’s approach to data privacy?
Q2. Were there any privacy strategies or features that reassured or concerned you?
Q3. What recommendations do you have for improving the system’s privacy features?
General Questions Q1. Which features of the system did you find most useful?
Q2. Were there any features that you did not use or found unnecessary? Why?
Q3. How do you think the system could be improved to better meet your needs?
Q4. Can you describe any specific scenarios where the system was particularly helpful?
Q5. How satisfied are you with the overall functionality of the system?
Q6. How do you think our system compares to other dietary monitoring tools in terms of advantages and disadvantages?
Q7. Is there anything else you would like to share about your experience with the system?

### v2

@@ 5.3.1 Evaluation Settings
Recognizing the variability in nutrient estimation, we recruited ten food and nutrition experts, including six nutrition interns, three researchers, and one Accredited Practicing Dietitian. They independently analyzed meal data (the ground truth meal images across the eating process), referencing the database provided by the Centre of Food Safety of the Government of the Hong Kong SAR database, with each meal randomly assigned to five experts for estimation. Each expert received about \$25 for compensation.
@@ 5.3.2 Evaluation Metrics
We measured expert agreement using the Intraclass Correlation Coefficient (ICC), showing high reliability of the average estimation from experts with ICC(2,k) > 0.9 (more details in Supplement 7). Therefore, to evaluate DietGlance’s performance, we compared its estimation to the expert averages using Mean Absolute Percentage Error (MAPE) due to varying nutrient units and scales, calculated as: ${MAPE=\frac{1}{N}\sum_{n=1}^{N}|\frac{\overline{E}_{n}-S_{n}}{\overline{E}_{n}}|\times 100\%}$, where $N$ is the number of meals. $\overline{E}_{n}$ and $S_{n}$ are the average estimations of experts and the system, respectively.
@@ 6 Study II: Four-Week Longitudinal Study
To evaluate the long-term effect of DietGlance on users’ dietary patterns, we recruited 16 participants from Study I for a 4-week longitudinal study with IRB approval, including 4 females and 12 males (P1, P6, P7, P10, P13, P15, P18, P20, P21, exp., P3, P8, P9, P14, P16, P17, P22, inexp.), aged 24 to 33 ($M=28.3$, $SD=2.4$). Building on the design requirements and findings from Study I, this study aims to explore how exp. and inexp. participants perceive and use DietGlance over time, and how it affects their dietary patterns.
@@ 6.1 Study Settings
  • Incorporated users’ habits (e.g., drinking Coke Zero instead of regular Coke) into prompts to refine diet identification.
@@
During the study, participants were encouraged to use Aria Glasses covering as many ingestive behaviors (meals and snacks) as possible to gain comprehensive insights into their habits and use the interface to review their dietary monitoring results. The interface was available 24/7, except from 12:00 am to 5:59 am for potential maintenance and backup, and was tracked via Google Analytics. Meal data older than seven days was archived, with analysis based on the most recent seven days to reflect current dietary patterns. Participants provided weekly feedback through questionnaires and interviews (details in Supplement 4). Participants received about \$17 every week and could review their monitoring and analysis results whenever the interface was available.
@@ 6.2.3 Long-Term General Usability, User Satisfaction, and User Experience
Usability improved throughout the study, with SUS scores rising from 76.09 in Study I to 85.00 by the end of Study II, reflecting increased perceived usability. Exp. participants consistently rated the system higher, while inexp. ones showed steady improvement, highlighting adaptability. NPS scores also rose, indicating sustained satisfaction across both groups. User experience, measured through UEQ dimensions, showed positive trends in attractiveness, efficiency, and dependability, while stimulation and novelty declined slightly, reflecting reduced novelty as users became familiar with the system. These findings emphasize the DietGlance’s high usability, satisfaction, and user experience, highlighting the need for tailoring onboarding processes and maintaining engagement, particularly for inexp. users. More details are in Supplement 8.
@@ 6.2.4 Long-Term Interaction Behaviors with DietGlance
Participants demonstrated consistent engagement with the system over four weeks, averaging 38.13 usage sessions. Exp. showed slightly higher engagement than inexp. ones. Initial novelty drove higher interactions in Week 1, which gradually stabilized into routine usage by Week 3. The average engagement time per session was 5.95 minutes, with exp. participants spending longer on average, reflecting deeper interaction with the system. The relatively low number of logged updates to diet summaries highlighted the system’s accuracy, while sustained chatbot interactions demonstrated its role in enhancing personalization and feedback. More details are available in the Supplement 9.
@@ 7 Discussion
In this section, we further discuss the implications of this work, followed by a discussion on the generalizability and limitations of this work.
7.1 Implications
@@ 7.3 Limitations and Future Work
The Aria Glasses, though powerful, are heavy due to unnecessary sensors for our study. Participants reported discomfort from sweat, especially in summer. Many also required vision correction, but lens replacement was not practical. This discomfort may hinder long-term retention. Future research could explore lighter glasses, attachable sensors for personal eyewear, or alternatives like smart earbuds or watches.
@@
DietGlance accurately identified most foods but struggled with visually similar items. While DietGlance accurately identified most foods, it struggled with visually similar items. Users may need to manually update diet summaries. The system also assumed equal portions in shared meals, potentially leading to discrepancies. Additionally, estimating micronutrients like phosphorus and potassium was challenging due to variability in biological, environmental, and processing factors. Future work could focus on improving micronutrient datasets for higher accuracy. As emphasized by the experts, higher precise nutrient intake was expected in clinical contexts.